OpenStudy (anonymous):

Final Balance~ $1317.28 APR(if needed)~ 5% Principle was $300, with monthly compounding for three years. Final interest total was 117.28 What percent of your final balance after the three years is due solely to interest? Now, suppose you invested for 30 years rather than 3, what is your final balance? Of that 30 yr balance, how much is interest alone? What percent of the 30 year balance is due soley to interest?

OpenStudy (anonymous):

Now, suppose you invested for 30 years rather than 3, what is your final balance? 300(1+.05/12)^(12*30) =1340.32 Of that 30 yr balance, how much is interest alone? 1317.28-117.28 =1200 1340.32-1200 =140.20

OpenStudy (anonymous):

What percent of your final balance after the three years is due solely to interest? 117.8/1317.28= 9% What percent of the 30 year balance is due solely to interest? 140.2/1340.32= 10.5%

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

You're welcome, hey do you know anything about trig functions and equations?

OpenStudy (anonymous):

A bit, yes.